# Saumili Janaa,1, John Kolinskib,2, Detlef Lohsea,c,3, and Vatsal Sanjaye,4

### v2

@@ 1 Introduction
Impacts of spherical bodies on rigid substrates span two classical limits that have long been treated separately: liquid drops wagner1932stoss, Philippi2016, zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025, sanjayUnifyingTheoryScaling2025 and elastic solids hertzUeberBeruhrungFester1882, pao1955extension, hunterEnergyAbsorbedElastic1957, bertinSimilaritySolutionsElastohydrodynamic2024. Both occur widely in nature and technology, where the normal force on the substrate is often the quantity of interest because it can damage engineered surfaces turbine_blade_wear2013, cheng2021drop. Drop impact governs processes from inkjet printing lohse2022fundamental and spray cooling/coating cooling_2017, spray_cool_Kim_2007 to forensics forensics_2018, pesticide deposition pesticide, and soil erosion Nearing1986. Impacts of elastic solids arise in hardness testing tirupataiah1991, granular media and suspensions granularmedia_Andreotti_2013, sports sports, and everyday bouncing of soft rubber balls. Despite this breadth, a unifying framework for the impact force across liquid and solid limits has remained elusive.
A falling liquid drop, after impact on a rigid surface deforms and spreads laterally until it reaches its maximum extent. A pronounced peak in the temporal evolution of force occurs at the instance of drop touchdown on the surface due to the inertia of the impact, whereas during droplet spreading this force is much smaller sanjayUnifyingTheoryScaling2025. For perfectly wetting surfaces, the liquid sticks to it. However, for non-wetting surfaces, the drop retracts from its maximum spread and generates a Worthington jet which coincides with a second peak in the temporal evolution of force zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025. In the inertial regime, Wagner’s theory predicts that the impact force scales as
@@
where $\rho_{l}$ is the density of the media, $V_{0}$ is the impact velocity and $R_{0}$ is the radius of the falling drop cheng2021drop.
By contrast, the elastic solids undergo slight deformation on impact with the substrate and bounce off following a brief contact with the substrate due to the exerted normal reaction. For such cases, the temporal evolution of force is characterized by a single maximum. Assuming the contact area to be small in comparison to the bead’s size, and considering a non-adhesive contact with small strains within the elastic limit, the situation can be treated as a Hertzian contact problem bertinSimilaritySolutionsElastohydrodynamic2024. Thus, Hertz’s theory describes the scaling laws for the impact force in the case of the impact of an elastic bead on a rigid substrate,
@@
Viscoelastic media – here, soft elastic gels – bridge liquids and solids: when deformed they support both viscous flow and recoverable elastic stress Snoeijer_etal_2020. Compared with Newtonian liquids, their rate‑dependent rheology can markedly alter spreading, pinch‑off, and rebound on impact clasen2006jfm, viscoelastic_drp_oscil_2021. Such soft media are relevant to inkjet printing inkjet_materials_Calvert, drop deposition control_deposition_visc, and spray atomization spray_atomisation_visc. Soft solids such as hydrogels, comprising cross‑linked networks with tunable elasticity, are widely used as biocompatible materials in rapid prototyping prototype_derby_science and drug delivery drug_delivery; see soft_materials_annrev for background.
@@ 2.1 Problem description and governing equations
The elastic stresses arise from deformation of the microstructure quantified by the conformation tension $\bm{\mathcal{A}}$ Snoeijer_etal_2020. Using the Oldroyd‑B constitutive model Oldroyd_1950,
@@
and Oldroyd‑B equivalent to a neo‑Hookean solid Snoeijer_etal_2020.
Despite being widespread due to simplicity, the Oldroyd-B model suffers from certain limitations Snoeijer_etal_2020, dixitViscoelasticWorthingtonJets2025. It fails to capture the shear-thinning behaviour in viscoelastic fluids completely, and erroneously predicts unbounded stress growth in strong extensional flows Mckinley_exten_flow. This limitation can be addressed by incorporating the finite polymer extension (FENE-P model)FENE-P_original. Also various other extensions of the Oldroyd-B model have been developed tanner2000 like the Phan-Thien-Tanner (PTT)Thien_1977 model to account for such non-linearities. However, since we are not dealing with strong extensional flows, we restrict ourselves to the Oldroyd-B model in this study, as it is sufficient to describe our case.
@@ 2.2 Numerical methods and simulations
We solve the above equations with Basilisk C popinet-basilisk, using a one‑fluid formulation with surface tension as a singular interfacial force tryggvason2011direct, brackbill1992continuum. The liquid–gas interface is tracked by a volume‑of‑fluid (VoF) color function $\psi$ advected by
@@
where the curvature $\kappa$ is computed via height functions popinet2018numerical. Explicit surface‑tension forcing imposes the standard capillary time‑step constraint popinet2009; the explicit update of $\bm{\sigma_{e}}$ adds a typically milder constraint.
At the substrate we impose no‑penetration and no‑slip, and a zero normal pressure gradient. To enforce a non‑contacting (superhydrophobic) condition we set $\psi=0$ at the wall, maintaining a thin air cushion kolinski-2014-epl, Vatsal_thesis. We stress that contact initiation in soft–solid impacts is generically air‑mediated and can proceed annularly or patchily with a non‑monotonic initial contact radius; see Zheng et al. (2021)zheng2021air for direct observations of air-mediated contact in compliant‑hemisphere impacts. Top and lateral boundaries use outflow (ambient pressure, zero tangential stress, zero normal velocity gradient). Boundaries are positioned far enough to avoid spurious confinement effects. The axisymmetric domain size is $8R_{0}\times 8R_{0}$. We employ quadtree adaptive mesh refinement (AMR) popinet2009, popinet2015 with maximal refinement at the interface and in regions of large velocity gradients. Wavelet‑based error control uses tolerances $10^{-3}$ for $\bm{u}$, $\psi$, $\kappa$, and $\bm{\mathcal{A}}$. Grid‑independence tests confirm convergence. Unless stated otherwise, the minimum cell size is $\Delta=R_{0}/512$ (i.e. 512 cells per radius on a uniform equivalent grid), increased to $\Delta=R_{0}/2048$ when required. Further numerical details can be found in popinet2015, Vatsal_thesis, dixitViscoelasticWorthingtonJets2025.
@@ 3 Wagner versus Hertz: Permanent‑Memory Impacts
We sweep the $(Ec,We)$ space over $We\in[1,10^{2}]$ and $Ec\in[10^{-1},10^{4}]$. The normal reaction on the substrate is computed from the drop’s momentum balance (eq. 12). For liquid drops on non-wetting substrates, $F(t)$ exhibits two peaks: an inertial peak at touchdown and a later peak associated with the formation of a Worthington jet zhangImpactForcesWater2022. The second, jetting peak occurs at a time $t_{2}\gg t_{\max}$, with the ratio $t_{2}/t_{\max}\sim\sqrt{We}$ sanjayRoleViscosityDrop2025, reflecting the phase difference between recoil and jet formation. In contrast, for elastic spheres the loading and unloading remain nearly in phase: the entire contact–rebound cycle fits within $t\lesssim 2t_{\max}$ and $F(t)$ displays only a single, almost symmetric peak in this interval. To compare liquids and solids and to track the transition, we therefore focus on the first (inertial) peak $F_{\max}$, non-dimensionalized as $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$; the corresponding time is $t_{\max}$.
Figure 2 shows representative cases across the parameter space. At low $Ec$ the sphere flows and behaves liquid‑like (figs. 2b,c; $Ec=1,2$). At low $Ec$ and low $We$, capillarity is significant (fig. 2b). At high $Ec$ the sphere deforms slightly and rebounds after a short contact (figs. 2d; $Ec=1000$). Increasing $We$ at fixed $Ec$ effectively softens the response (reduces $El$) and increases the contact duration (figs. 2d). The force traces reflect this evolution: for large $El$ (high $Ec$, low $We$) $F(t)$ is nearly symmetric, as in elastic impacts, while decreasing $El$ (e.g. by increasing $We$) skews $F(t)$ in the manner typical of liquid impacts zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025. The peak magnitude also varies appreciably across cases.
Fig. 2: (a) Phase space in the EcEc-WeWe plane illustrating the range of simulations conducted in this work colored according to the elasticity number $El=Ec/We$. The four highlighted symbols locate typical cases representing the range of parameters explored. We chose ($We,Ec$) == (b) ($5,1$), (c) ($500,2$), (d) ($5,1000$), (e) ($500,1000$). For each case, the color scheme of each snapshot represents the magnitude of the velocity normalized by the impact velocity, alongside the corresponding force history $F(t)/(\rho_{l}V_{0}^{2}R_{0}^{2})$ plotted versus $t/t_{\max}$ (right). The force traces are plotted up to $t/t_{\max}=2$: for the liquid-drop reference, the second peak associated with the Worthington jet zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025 occurs at later times $t_{2}\gg t_{\max}$ (with $t_{2}/t_{\max}\sim\sqrt{We}$) and is therefore outside the plotted window.
The dependence of $F_{\max}$ on $We$ and $Ec$ is summarized in figure 3. For $Ec\lesssim 1$, $F_{\max}$ follows the liquid‑impact trend with the low‑$We$ correction zhangImpactForcesWater2022,
@@
Collapsing the data using $El$ (figure 4a) reveals two regimes. For $El\lesssim 1$ and sufficiently large $We$, the data lie near Wagner’s constant level, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$. At small $We$, inertia competes with capillarity and the low‑$We$ correction in eq. (21) is required Vatsal_thesis. For $El\gtrsim 1$, all points collapse onto a single master curve with slope $2/5$:
@@ 4.1 Purely Elastic Limit: Hertz Contact Theory
Consider a solid elastic sphere of radius $R_{0}$, mass $m$, and elastic modulus $G$ (shear modulus, assuming an incompressible material) impacting a rigid flat surface with speed $V_{0}$. Upon contact, the sphere deforms and a normal force $F$ develops according to Hertz’s contact law. For a sphere indenting a half-space, the force–indentation relation is given by the 3/2-power law of classical Hertz contact mechanics hertzUeberBeruhrungFester1882,
@@ 4.2 Purely Liquid Limit: Wagner Impact Theory
At the opposite extreme ($Ec\to 0$) the sphere behaves as a liquid drop, and its impact dynamics are governed by inertia and capillarity in the classical Wagner limit cheng2021drop, zhangImpactForcesWater2022. Instead of an elastic compression, the drop undergoes rapid localized deformation at the moment of impact: the south pole flattens against the substrate while the remainder of the drop (including the north pole) is still moving downward at nearly the impact speed. The vertical momentum of the drop’s center of mass is redirected into a radial outflow along the substrate, causing a small “wetted” area to grow outward from the impact pointsanjayRoleViscosityDrop2025. This Wagner-type mechanism – a thin spreading lamella initiated at the contact point Eggers2010, Josserand2016, negusModellingDropletImpact2022, negus2021droplet – contrasts sharply with the distributed Hertzian contact of an elastic solid. It produces a pronounced impulsive force at touchdown, as the drop’s momentum is arrested over a short time and small area. The first force peak thus originates from pure inertial impingement of the liquid on the surfacesanjayUnifyingTheoryScaling2025. We have analyzed this case in past, for details see the references zhangImpactForcesWater2022, sanjayRoleViscosityDrop2025, sanjayUnifyingTheoryScaling2025.
@@
During the very early stage ($t\sim\tau_{\rho}=D_{0}/V_{0}$), the normal force rises sharply to its first maximum $F_{\max}$ as the drop’s inertia is transferred to the substrate. At this moment the deformation is still localized: the contact radius has grown only to the order of the drop’s initial radius. In fact, experiments confirm that at the peak force time $t_{\max}$, the spread diameter $D_{f}(t_{\max})$ is approximately equal to the initial drop diameter $D_{0}$, consistent with early-time self-similarity of the impact sanjayRoleViscosityDrop2025, sanjayUnifyingTheoryScaling2025, Mandre_etalPRL_2019, Philippi_Lagrée_Antkowiak_2016. Wagner’s inviscid theory predicts that the peak force scales with the inertial pressure on the drop’s footprint. Non-dimensionalizing $F_{\text{max}}$ by $\rho_{l}V_{0}^{2}R_{0}^{2}$ (with $\rho_{l}$ the liquid density) yields a constant of order unity. Indeed, for large Weber numbers (negligible surface tension), simulations and experiments find $F_{\text{max}}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$, cf. figure 3a of zhangImpactForcesWater2022. At lower $We$, surface tension enhances the impact, and the peak force is increases with decreasing $We$ (following a $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim We^{-1}$ correction in this regime). This initial peak is inertia-dominated and is relatively insensitive to liquid viscosity: $F_{\text{max}}$ remains nearly constant for drops with viscosity up to about $100\times$ that of water sanjayRoleViscosityDrop2025. Since our spheres have a very low background viscosity ($Oh=10^{-2}$), in the limit $Ec\to 0$ our results for $F_{\text{max}}$ follow the same trends. Only for highly viscous drops (Ohnesorge number $Oh\gtrsim 1$) does viscous dissipation significantly attenuate the first peak, reflecting the fact that most of the drop’s momentum is redirected (and the force generated) before substantial viscous effects have time to act, cf. sanjayUnifyingTheoryScaling2025.

### v3

@@ 1 Introduction
Impacts of spherical bodies on rigid substrates span two classical limits that have long been treated separately: liquid drops 74, 48, 76, 56, 1 and elastic solids 8, 47, 23, 30. Both occur widely in nature and technology, where the normal force on the substrate is often the quantity of interest because it can damage engineered surfaces 3, 12. Drop impact governs processes from inkjet printing 36 and spray cooling/coating 59, 27 to forensics 60, pesticide deposition 22, and soil erosion 43. Impacts of elastic solids arise in hardness testing 71, granular media and suspensions 4, sports 16, and everyday bouncing of soft rubber balls. Despite this breadth, a unifying framework for the impact force across liquid and solid limits has remained elusive.
A falling liquid drop, after impact on a rigid surface deforms and spreads laterally until it reaches its maximum extent. A pronounced peak in the temporal evolution of force occurs at the instance of drop touchdown on the surface due to the inertia of the impact, whereas during droplet spreading this force is much smaller 1. For perfectly wetting surfaces, the liquid sticks to it. However, for non-wetting surfaces, the drop retracts from its maximum spread and generates a Worthington jet which coincides with a second peak in the temporal evolution of force 76, 56. In the inertial regime, Wagner’s theory predicts that the impact force scales as
@@
where $\rho_{l}$ is the density of the media, $V_{0}$ is the impact velocity and $R_{0}$ is the radius of the falling drop 12.
By contrast, the elastic solids undergo slight deformation on impact with the substrate and bounce off following a brief contact with the substrate due to the exerted normal reaction. For such cases, the temporal evolution of force is characterized by a single maximum. Assuming the contact area to be small in comparison to the bead’s size, and considering a non-adhesive contact with small strains within the elastic limit, the situation can be treated as a Hertzian contact problem 30. Thus, Hertz’s theory describes the scaling laws for the impact force in the case of the impact of an elastic bead on a rigid substrate,
@@
Viscoelastic media – here, soft elastic gels – bridge liquids and solids: when deformed they support both viscous flow and recoverable elastic stress 63. Compared with Newtonian liquids, their rate‑dependent rheology can markedly alter spreading, pinch‑off, and rebound on impact 15, 65. Such soft media are relevant to inkjet printing 9, drop deposition 5, and spray atomization 68. Soft solids such as hydrogels, comprising cross‑linked networks with tunable elasticity, are widely used as biocompatible materials in rapid prototyping 17 and drug delivery 35; see Chen et al. 11 for background.
@@ 2.1 Problem description and governing equations
The elastic stresses arise from deformation of the microstructure quantified by the conformation tension $\bm{\mathcal{A}}$ 63. Using the Oldroyd‑B constitutive model 46,
@@
and Oldroyd‑B equivalent to a neo‑Hookean solid 63.
Despite being widespread due to simplicity, the Oldroyd-B model suffers from certain limitations 63, 18. It fails to capture the shear-thinning behaviour in viscoelastic fluids completely, and erroneously predicts unbounded stress growth in strong extensional flows 40. This limitation can be addressed by incorporating the finite polymer extension (FENE-P model)6. Also various other extensions of the Oldroyd-B model have been developed 66 like the Phan-Thien-Tanner (PTT)67 model to account for such non-linearities. However, since we are not dealing with strong extensional flows, we restrict ourselves to the Oldroyd-B model in this study, as it is sufficient to describe our case.
@@ 2.2 Numerical methods and simulations
We solve the above equations with Basilisk C 50, using a one‑fluid formulation with surface tension as a singular interfacial force 72, 7. The liquid–gas interface is tracked by a volume‑of‑fluid (VoF) color function $\psi$ advected by
@@
where the curvature $\kappa$ is computed via height functions 51. Explicit surface‑tension forcing imposes the standard capillary time‑step constraint 52; the explicit update of $\bm{\sigma_{e}}$ adds a typically milder constraint.
At the substrate we impose no‑penetration and no‑slip, and a zero normal pressure gradient. To enforce a non‑contacting (superhydrophobic) condition we set $\psi=0$ at the wall, maintaining a thin air cushion 28, 57. For liquid drops, enforcing this air cushion results in a non-wetting (superhydrophobic) substrate, while for elastic spheres this results in a non-contacting substrate. We stress that contact initiation in soft–solid impacts is generically air‑mediated and can proceed annularly or patchily with a non‑monotonic initial contact radius; see Zheng et al. (2021)78 for direct observations of air-mediated contact in compliant‑hemisphere impacts. Top and lateral boundaries use outflow (ambient pressure, zero tangential stress, zero normal velocity gradient). Boundaries are positioned far enough to avoid spurious confinement effects. The axisymmetric domain size is $8R_{0}\times 8R_{0}$. We employ quadtree adaptive mesh refinement (AMR) 52, 53 with maximal refinement at the interface and in regions of large velocity gradients. Wavelet‑based error control uses tolerances $10^{-3}$ for $\bm{u}$, $\psi$, $\kappa$, and $\bm{\mathcal{A}}$. Grid‑independence tests confirm convergence. Unless stated otherwise, the minimum cell size is $\Delta=R_{0}/512$ (i.e. 512 cells per radius on a uniform equivalent grid), increased to $\Delta=R_{0}/2048$ when required. Further numerical details can be found in Popinet 53, Sanjay 57 and Dixit et al. 18.
@@ 3 Wagner versus Hertz: Permanent‑Memory Impacts
We sweep the $(Ec,We)$ space over $We\in[1,10^{3}]$ and $Ec\in[10^{-1},10^{4}]$. The normal reaction on the substrate is computed from the drop’s momentum balance (eq. 12). For liquid drops on non-wetting substrates, $F(t)$ exhibits two peaks: an inertial peak at touchdown and a later peak associated with the formation of a Worthington jet 76. The second, jetting peak occurs at a time $t_{2}\gg t_{\max}$, with the ratio $t_{2}/t_{\max}\sim\sqrt{We}$ 56, reflecting the phase difference between recoil and jet formation. In contrast, for elastic spheres the loading and unloading remain nearly in phase: the entire contact–rebound cycle fits within $t\lesssim 2t_{\max}$ and $F(t)$ displays only a single, almost symmetric peak in this interval. To compare liquids and solids and to track the transition, we therefore focus on the first (inertial) peak $F_{\max}$, non-dimensionalized as $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})$; the corresponding time is $t_{\max}$.
Fig. 2: (a) Phase space in the EcEc-WeWe plane illustrating the range of simulations conducted in this work colored according to the elasticity number $El=Ec/We$. The four highlighted symbols locate typical cases representing the range of parameters explored. We chose ($We,Ec$) == (b) ($5,1$), (c) ($500,2$), (d) ($5,1000$), (e) ($500,1000$). For each case, the color scheme of each snapshot represents the magnitude of the velocity normalized by the impact velocity, alongside the corresponding force history $F(t)/(\rho_{l}V_{0}^{2}R_{0}^{2})$ plotted versus $t/t_{\max}$ (right). The force traces are plotted up to $t/t_{\max}=2$: for the liquid-drop reference, the second peak associated with the Worthington jet 76, 56 occurs at later times $t_{2}\gg t_{\max}$ (with $t_{2}/t_{\max}\sim\sqrt{We}$) and is therefore outside the plotted window.
Figure 2 shows representative cases across the parameter space. At low $Ec$ the sphere flows and behaves liquid‑like (figs. 2b,c; $Ec=1,2$). At low $Ec$ and low $We$, capillarity is significant (fig. 2b). At high $Ec$ the sphere deforms slightly and rebounds after a short contact (figs. 2d; $Ec=1000$). Increasing $We$ at fixed $Ec$ effectively softens the response (reduces $El$) and increases the contact duration (figs. 2d). The force traces reflect this evolution: for large $El$ (high $Ec$, low $We$) $F(t)$ is nearly symmetric, as in elastic impacts, while decreasing $El$ (e.g. by increasing $We$) skews $F(t)$ in the manner typical of liquid impacts 76, 56. The peak magnitude also varies appreciably across cases. Since force and time are related as $F_{max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\cdot t_{max}/(R_{0}/V_{0})\sim 1$ for any impacting sphere in general, $t_{max}$ reduces accordingly with an increase in $F_{max}$.
The dependence of $F_{\max}$ on $We$ and $Ec$ is summarized in figure 3. For $Ec\lesssim 1$, $F_{\max}$ follows the liquid‑impact trend with the low‑$We$ correction 76,
@@
Collapsing the data using $El$ (figure 4a) reveals two regimes. For $El\lesssim 1$ and sufficiently large $We$, the data lie near Wagner’s constant level, $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$. At small $We$, inertia competes with capillarity and the low‑$We$ correction in eq. (21) is required 57. For $El\gtrsim 1$, all points collapse onto a single master curve with slope $2/5$:
@@ 4.1 Purely Elastic Limit: Hertz Contact Theory
Consider a solid elastic sphere of radius $R_{0}$, mass $m$, and elastic modulus $G$ (shear modulus, assuming an incompressible material) impacting a rigid flat surface with speed $V_{0}$. Upon contact, the sphere deforms and a normal force $F$ develops according to Hertz’s contact law. For a sphere indenting a half-space, the force–indentation relation is given by the 3/2-power law of classical Hertz contact mechanics 8,
@@ 4.2 Purely Liquid Limit: Wagner Impact Theory
At the opposite extreme ($Ec\to 0$) the sphere behaves as a liquid drop, and its impact dynamics are governed by inertia and capillarity in the classical Wagner limit 12, 76. Instead of an elastic compression, the drop undergoes rapid localized deformation at the moment of impact: the south pole flattens against the substrate while the remainder of the drop (including the north pole) is still moving downward at nearly the impact speed. The vertical momentum of the drop’s center of mass is redirected into a radial outflow along the substrate, causing a small “wetted” area to grow outward from the impact point56. This Wagner-type mechanism – a thin spreading lamella initiated at the contact point 20, 26, 45, 44 – contrasts sharply with the distributed Hertzian contact of an elastic solid. It produces a pronounced impulsive force at touchdown, as the drop’s momentum is arrested over a short time and small area. The first force peak thus originates from pure inertial impingement of the liquid on the surface1. We have analyzed this case in past, for details see the references Zhang et al. 76, Sanjay et al. 56, Sanjay and Lohse 1.
@@
During the very early stage ($t\sim\tau_{\rho}=D_{0}/V_{0}$), the normal force rises sharply to its first maximum $F_{\max}$ as the drop’s inertia is transferred to the substrate. At this moment the deformation is still localized: the contact radius has grown only to the order of the drop’s initial radius. In fact, experiments confirm that at the peak force time $t_{\max}$, the spread diameter $D_{f}(t_{\max})$ is approximately equal to the initial drop diameter $D_{0}$, consistent with early-time self-similarity of the impact 56, 1, 38, Philippi_Lagrée_Antkowiak_2016. Wagner’s inviscid theory predicts that the peak force scales with the inertial pressure on the drop’s footprint. Non-dimensionalizing $F_{\text{max}}$ by $\rho_{l}V_{0}^{2}R_{0}^{2}$ (with $\rho_{l}$ the liquid density) yields a constant of order unity. Indeed, for large Weber numbers (negligible surface tension), simulations and experiments find $F_{\text{max}}/(\rho_{l}V_{0}^{2}R_{0}^{2})\approx 3.24$, cf. figure 3a of Zhang et al. 76. At lower $We$, surface tension enhances the impact, and the peak force is increases with decreasing $We$ (following a $F_{\max}/(\rho_{l}V_{0}^{2}R_{0}^{2})\sim We^{-1}$ correction in this regime). This initial peak is inertia-dominated and is relatively insensitive to liquid viscosity: $F_{\text{max}}$ remains nearly constant for drops with viscosity up to about $100\times$ that of water 56. Since our spheres have a very low background viscosity ($Oh=10^{-2}$), in the limit $Ec\to 0$ our results for $F_{\text{max}}$ follow the same trends. Only for highly viscous drops (Ohnesorge number $Oh\gtrsim 1$) does viscous dissipation significantly attenuate the first peak, reflecting the fact that most of the drop’s momentum is redirected (and the force generated) before substantial viscous effects have time to act, cf. Sanjay and Lohse 1.